<Homework #2 Problem 20.60 MC 6 of7 > Corect Review 1 Constants Significant Figures Feedback: Your answer 5.76-10-1576x013 N was either rounded differently or used a different number of significant figures than required for this part Two 2.0-cm-diameter disks are parallel to each other, 2.0 mm apart. They are charged to ± 10 nC Part A proton is shot from the negative disk toward the positive disk. What launch speed must the proton have to just barely reach the positive disk? Assume the weight of the proton is negligible Express your answer using three significant figures. View Available Hints 지 X. 101
Answer #1

Electric field, E = Q/pi r^2 e0

E = (10 x 10^-9)/(pi x 8.85 x 10^-12 x (0.01)^2 = 3.60 x 10^6 N/C

Acceleration, a = eE/m = (1.6 x 10^-19 x 3.6 x 10^6)/(1.67 x 10^-27) = 3.45 x 10^14 m/s^2

Time taken, t = sqrt(2 x 0.002/3.45 x 10^14) = 3.41 x 10^-9 sec

Velocity, v = (3.45 x 10^14)(3.41 x 10^-9) = 1.17 x 10^6 m/s

Comment in case any doubt please rate my answer ....